OpenStudy (liizzyliizz):

doman and range of y= 1/x

OpenStudy (anonymous):

domain all real x which does not equal to 0 range all reals except for zero

jimthompson5910 (jim_thompson5910):

You cannot divide by zero. So the domain is the set of all real numbers but x cannot equal zero. Because the horizontal asymptote is y = 0, this means that the range is the set of all real numbers but y cannot equal zero.

OpenStudy (liizzyliizz):

i see what u are saying but how would u write that in interval notation which is my biggest issue..

jimthompson5910 (jim_thompson5910):

Domain: \[(-\infty,0)\cup(0,\infty)\] Range: \[(-\infty,0)\cup(0,\infty)\] Note: in this case, the domain and range are the same. This is not always the case though.